OpenStudy (anonymous):

Find all the zeros of the equation x^3+x^2-3x+3=0

OpenStudy (tkhunny):

The coefficients are ripe for GROUPING!! x^3+x^2-3x+3=0 Consider the first two terms and then the last two terms. (x^2)(x+1) - 3(x+1) = 0 (x^2 - 3)(x+1) = 0 Where does that leave us?

OpenStudy (anonymous):

so what is the answer?

OpenStudy (tkhunny):

There is a reason I did not work the whole problem. You must assert your knowledge and do the rest. Prove that you understand the problem by completing the problem.